### v1

@@ 4.2 A Smooth Path to a Normal-Form Perfect Equilibrium
By employing the optimality conditions to the problem (13), we derive a polynomial system that coincides with the system (10). Hence, for any $(\gamma(t),t)\in\widetilde{\mathscr{F}}^{c}$, there exists a unique pair $(\lambda,\nu)$ such that the system (10) is satisfied. Let $\widetilde{\mathscr{C}}^{c}_{L}=\{(\gamma(t),t,\lambda,\nu)\in\widetilde{\mathscr{C}}_{L}|(\gamma(t),t)\in\widetilde{\mathscr{F}}^{c}\}$ and $\mathscr{C}^{c}_{L}$ be the closure of $\widetilde{\mathscr{C}}^{c}_{L}$. From the preceding discussion, we obtain that $\mathscr{C}^{c}_{L}$ forms a connected component within $\mathscr{C}_{L}$ that intersects $\mathbb{R}^{n_{0}}\times\{2\}\times\mathbb{R}^{n_{0}}\times\mathbb{R}^{m_{0}}$. Consider a convergent sequence $\{(\gamma(t_{k}),t_{k})\}_{k=1}^{\infty}\subseteq\widetilde{\mathscr{F}}^{c}$, where $\lim_{k\to\infty}t_{k}=0$. We associate each $(\gamma(t_{k}),t_{k})$ with the corresponding pair $(\lambda(t_{k}),\mu(t_{k}))$ such that $(\gamma(t_{k}),t_{k},\lambda(t_{k}),\mu(t_{k}))\in\widetilde{\mathscr{C}}^{c}_{L}$. The boundedness of $\widetilde{\mathscr{C}}^{c}_{L}$ guarantees the existence of a convergent subsequence of $\{(\gamma(t_{k}),t_{k},\lambda(t_{k}),\mu(t_{k}))\}_{k=1}^{\infty}$. As a result, $\mathscr{C}^{c}_{L}$ intersects $\mathbb{R}^{n_{0}}\times\{0\}\times\mathbb{R}^{n_{0}}\times\mathbb{R}^{m_{0}}$, thereby completing the proof. ∎
@@
Lemma 6 asserts that the connected component discussed in Theorem 5 is uniquely determined and intersects the level $t=2$ at the point $(\gamma^{*}(2),2,\lambda^{*}(2),\nu^{*}(2))$. In order to eliminate the impact of $\lambda^{i}(w^{i}_{I^{j}_{i}}a)=0$ for $i\in N,(j,a)\notin D_{i}$ on the differentiability of the path over the interval $(0,1]$, and reduce the number of variables for more efficient computation, we employ a variable substitution technique as outlined in Cao et al. [44]. Given $\tau_{0}>0$ and $\kappa_{0}>1$, define the following functions,
@@
It follows that $\psi_{1}(v,r;\tau_{0},\kappa_{0})\psi_{2}(v,r;\tau_{0},\kappa_{0})=(\tau_{0}r)^{\kappa_{0}}$. Since $\kappa_{0}>1$, both functions are continuously differentiable on the domain ×$\mathbb{R}\times(0,\infty)$. For $x=(x^{i}(w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))\in\mathbb{R}^{n_{0}}$, we define $\gamma(x,t)=(\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$ and $\lambda(x,t)=(\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$, where
$\begin{array}[]{l}\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\left\{\begin{array}[]{ll}\rho(t)(1-\theta(t))\eta^{0i}(w^{i}_{I^{j}_{i}}a)+\psi_{1}(x^{i}(w^{i}_{I^{j}_{i}}a),c(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})&(j,a)\in D_{i},\\ \psi_{1}(x^{i}(w^{i}_{I^{j}_{i}}a),\theta(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})&(j,a)\notin D_{i},\end{array}\right.\\ \lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\left\{\begin{array}[]{ll}\psi_{2}(x^{i}(w^{i}_{I^{j}_{i}}a),c(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})&(j,a)\in D_{i},\\ \psi_{2}(x^{i}(w^{i}_{I^{j}_{i}}a),\theta(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})&(j,a)\notin D_{i}.\end{array}\right.\end{array}$ (15)
It is evident that $(\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)-\rho(t)(1-\theta(t))\eta^{0i}(w^{i}_{I^{j}_{i}}a))\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)=c(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a)$ for $i\in N,(j,a)\in D_{i}$ and $\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a)$ for $i\in N,(j,a)\notin D_{i}$. Let $\alpha=(\alpha(w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))\in\mathbb{R}^{n_{0}}$ be an arbitrary vector with $||\alpha||$ sufficiently small. By substituting $\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)$ and $\lambda^{i}(w^{i}_{I^{j}_{i}}a)$ by $\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)$ and $\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)$ in the system (10) and subtracting $c(t)(1-\theta(t))\alpha$ from the left-hand side in the first group of equations, we obtain an equivalent formulation with fewer variables and constraints,
$\begin{array}[]{l}(1-c(t))g^{i}(w^{i}_{I^{j}_{i}}a,\gamma^{-i}(x,t))+\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)\\ \hskip 113.81102pt-c(t)(1-\theta(t))\alpha(w^{i}_{I^{j}_{i}}a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(x,t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,i\in N,j\in M_{i}.\\ \end{array}$ (16)
@@
Given almost any $\alpha\in\mathbb{R}^{n_{0}}$ with sufficiently small $||\alpha||$, a smooth path can be identified in $\mathscr{P}_{L}$ that begins at $(x^{*}(2),2,\nu^{*}(2))$ when $t=2$ and leads to a normal-form perfect equilibrium as $t$ approaches zero.
@@
Let $p(x,t,\nu;\alpha)$ denote the left-hand sides of the equations in the system (16), and define $p_{\alpha}(x,t,\nu)=p(x,t,\nu;\alpha)$ when $\alpha$ is treated as a constant. The function $p(x,t,\nu;\alpha)$ is continuously differentiable over $\mathbb{R}^{n_{0}}\times(0,2)\times\mathbb{R}^{m_{0}}\times\mathbb{R}^{n_{0}}$. As demonstrated in Appendix B, its Jacobian matrix has full-row rank in this region. Using the transversality theorem as stated by Eaves and Schmedders [45], it can be shown that zero is a regular value of $p_{\alpha}(x,t,\nu)$ over $\mathbb{R}^{n_{0}}\times(0,2)\times\mathbb{R}^{m_{0}}$ for almost any $\alpha$ with $||\alpha||<\epsilon$, where $\epsilon$ is a sufficiently small positive constant.
We choose a suitable $\alpha$ such that zero is a regular value of $p_{\alpha}(x,t,\nu)$ over $\mathbb{R}^{n_{0}}\times(0,2)\times\mathbb{R}^{m_{0}}$. By applying the implicit function theorem, it can be inferred that the component described in Theorem 5 derives a smooth path in $\mathscr{P}_{L}$ that initiates at $((x^{*}(2),2,\nu^{*}(2))$ and terminates at $t=0$. In Appendix B, we show that, at $t=2$, zero remains a regular value of $p_{\alpha}(x,2,\nu)$ in $\mathbb{R}^{n_{0}}\times\mathbb{R}^{m_{0}}$, ensuring the path does not intersect tangentially with $\mathbb{R}^{n_{0}}\times\{2\}\times\mathbb{R}^{m_{0}}$. Theorems 3 then confirm that this smooth path ultimately yields a normal-form perfect equilibrium. This completes the proof. ∎
As the third summation term in the optimization objective of (9) vanishes when $t\in(0,1]$, we modify the variable substitution (15) for $(j,a)\notin D_{i}$ to achieve a smoother transition at $t=1$. Specifically, we introduce the following adjustments,
$\begin{array}[]{l}\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)=(1-\theta(t))x^{i}(w^{i}_{I^{j}_{i}}a)+\theta(t)\psi_{1}(x^{i}(w^{i}_{I^{j}_{i}}a),c(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0}),\\ \lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\theta(t)\psi_{2}(x^{i}(w^{i}_{I^{j}_{i}}a),c(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0}),i\in N,(j,a)\notin D_{i}.\end{array}$ (17)
Under this substitution, the resulting system is equivalent to (16) for $t\in(0,1]\cup\{2\}$. Consequently, we obtain a smoother path that converges to a normal-form perfect equilibrium.
@@ 5 Extension of Harsanyi’s Tracing Procedures to the Sequence Form
According to the analysis in [46], the tracing procedure proposed by von Stengel et al. [36] for two-player extensive-form games is equivalent to Harsanyi’s linear tracing procedure when applied in sequence form. In subsection 5.1, we extend this tracing procedure to accommodate $n$-player games and develop an alternative differentiable path-following method for computing normal form perfect equilibria. Furthermore, to improve the smoothness of the equilibrium paths, we extend Harsanyi’s logarithmic tracing procedure to the sequence form in subsection 5.2, providing an additional differentiable path-following method.
@@ 5.1 Harsanyi’s Linear Tracing Procedure in Sequence Form
Let $p^{0}=(p^{0i}(w^{i}):i\in N,w^{i}\in W^{i})$ be a given prior realization plan profile. For $\gamma\in\Lambda$ and $t\in[0,1]$, define $y(\gamma,t)=(y(\gamma^{i},t;w^{i}):i\in N,w^{i}\in W^{i})$ with $y(\gamma^{i},t;w^{i})=(1-t)\gamma^{i}(w^{i})+tp^{0i}(w^{i})$. This construction ensures $y(\gamma,t)\in\Lambda$. For $t\in(0,1]$, we introduce an artificial game $\Gamma_{s}^{v}(t)$ in sequence form, where each player $i$ finds their best response to $\hat{\gamma}\in\Lambda$ by solving the following linear optimization problem,
$\begin{array}[]{rl}\max\limits_{\gamma^{i}}&\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(w^{i}_{I^{j}_{i}}a)g^{i}(w^{i}_{I^{j}_{i}}a,y(\hat{\gamma}^{-i},t))\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(w^{i}_{I^{j}_{i}}a)-\gamma^{i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ &0\leq\gamma^{i}(w^{i}_{I^{j}_{i}}a),\;(j,a)\in D_{i}.\end{array}$ (18)
@@
$\begin{array}[]{l}g^{i}(w^{i}_{I^{j}_{i}}a,y(\gamma^{-i},t))+\lambda^{i}(w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=0,\;i\in N,(j,a)\in D_{i},\\ g^{i}(w^{i}_{I^{j}_{i}}a,y(\gamma^{-i},t))-\nu^{i}_{I^{j}_{i}}+\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,(j,a)\notin D_{i},\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(w^{i}_{I^{j}_{i}}a)-\gamma^{i}(w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ \gamma^{i}(w^{i}_{I^{j}_{i}}a)\lambda^{i}(w^{i}_{I^{j}_{i}}a)=0,\;0\leq\gamma^{i}(w^{i}_{I^{j}_{i}}a),\;0\leq\lambda^{i}(w^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\in D_{i},\end{array}$ (19)
where $\zeta^{i}_{I^{j}_{i}}(a)=\sum_{j_{q}\in M_{i}(w^{i}_{I^{j}_{i}}a)}\nu^{i}_{I^{j_{q}}_{i}}$. As a result, the solution $\gamma^{*}$ solves the optimization problem (18) against itself if and only if a pair $(\lambda^{*},\nu^{*})$ exists together with $\gamma^{*}$ that collectively satisfies the system (19). Referring to the proof of Lemma 4, we deduce that the artificial game $\Gamma_{s}^{v}(t)$ can be equivalently reformulated as the perturbed game $\Gamma_{s}(t)$ with the perturbation $\eta(t)=(\eta^{i}(t;w^{i}):i\in N,w^{i}\in W^{i})$ defined by $\eta^{i}(t;w^{i})=tp^{0i}(w^{i})$. The strategy profile $\gamma^{*}$ is a Nash equilibrium of $\Gamma_{s}^{v}(t)$ whenever $y(\gamma^{*},t)$ constitutes a Nash equilibrium in $\Gamma_{s}(t)$. This leads to the following theorem.
@@
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(t)}&(1-\theta(t))\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)g^{i}(w^{i}_{I^{j}_{i}}a,y(\hat{\gamma}^{-i},\rho(t)))\\ &-\theta(t)\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{0i}(w^{i}_{I^{j}_{i}}a)\ln\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ &0\leq\gamma^{i}(t;w^{i}_{I^{j}_{i}}a),\;(j,a)\in D_{i}.\end{array}$ (20)
@@
The application of the optimality conditions to the problem (20), coupled with the enforcement $\hat{\gamma}=\gamma(t)$, results in a polynomial equilibrium system. This system is specified by (21) for $t\in(1,2]$ and is equivalent to (19) for $t\in(0,1]$.
@@
$\begin{array}[]{l}(1-\theta(t))g^{i}(w^{i}_{I^{j}_{i}}a,y(\gamma^{-i}(t),\rho(t)))+\lambda^{i}(w^{i}_{I^{j}_{i}}a)\\ \hskip 105.2751pt-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ \gamma^{i}(t;w^{i}_{I^{j}_{i}}a)\lambda^{i}(w^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a),\;0<\gamma^{i}(t;w^{i}_{I^{j}_{i}}a),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (21)

### v3

@@ 4.2 A Smooth Path to a Normal-Form Perfect Equilibrium
By applying the optimality conditions to the problem (13), we obtain a polynomial system equivalent to the system (10). Consequently, for each $(\gamma(t),t)\in\widetilde{\mathscr{F}}^{c}$, there exists a unique pair $(\lambda,\nu)$ such that the system (10) holds. Define $\widetilde{\mathscr{C}}^{c}_{L}=\{(\gamma(t),t,\lambda,\nu)\in\widetilde{\mathscr{C}}_{L}|(\gamma(t),t)\in\widetilde{\mathscr{F}}^{c}\}$ and let $\mathscr{C}^{c}_{L}$ be its closure. From the previous arguments, it follows that $\mathscr{C}^{c}_{L}$ constitutes a connected component within $\mathscr{C}_{L}$ intersecting $\mathbb{R}^{n_{0}}\times\{2\}\times\mathbb{R}^{n_{0}}\times\mathbb{R}^{m_{0}}$. Let $\{(\gamma(t_{k}),t_{k})\}_{k=1}^{\infty}\subseteq\widetilde{\mathscr{F}}^{c}$ be a sequence converging with $\lim_{k\to\infty}t_{k}=0$. For each element $(\gamma(t_{k}),t_{k})$, we associate a corresponding pair $(\lambda(t_{k}),\mu(t_{k}))$ such that $(\gamma(t_{k}),t_{k},\lambda(t_{k}),\mu(t_{k}))\in\widetilde{\mathscr{C}}^{c}_{L}$. Since $\widetilde{\mathscr{C}}^{c}_{L}$ is bounded, there exists a convergent subsequence of $\{(\gamma(t_{k}),t_{k},\lambda(t_{k}),\mu(t_{k}))\}_{k=1}^{\infty}$. Therefore, the set $\mathscr{C}^{c}_{L}$ intersects $\mathbb{R}^{n_{0}}\times\{0\}\times\mathbb{R}^{n_{0}}\times\mathbb{R}^{m_{0}}$, which concludes the proof. ∎
@@
Lemma 6 asserts that the connected component discussed in Theorem 5 is uniquely determined and intersects the level $t=2$ at the point $(\gamma^{*}(2),2,\lambda^{*}(2),\nu^{*}(2))$. In order to eliminate the impact of $\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=0$ for $i\in N,(j,a)\notin D_{i}$ on the differentiability of the path over the interval $(0,1]$, and reduce the number of variables for more efficient computation, we employ a variable substitution technique as outlined in Cao et al. [Caodifferentiablepathfollowingmethod2022]. Given $\tau_{0}>0$ and $\kappa_{0}>1$, define the following functions,
@@
It can be deduced that $\psi_{1}(v,r;\tau_{0},\kappa_{0})\psi_{2}(v,r;\tau_{0},\kappa_{0})=(\tau_{0}r)^{\kappa_{0}}$. Since $\kappa_{0}>1$, both functions possess continuous differentiability over the domain $\mathbb{R}\times(0,\infty)$. For $x=(x^{i}(\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))\in\mathbb{R}^{n_{0}}$, we define $\gamma(x,t)=(\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$ and $\lambda(x,t)=(\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$, where
[…]
It is evident that $(\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)-\rho(t)(1-\theta(t))\eta^{0i}(\varpi^{i}_{I^{j}_{i}}a))\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)=c(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)$ for $i\in N,(j,a)\in D_{i}$ and $\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)$ for $i\in N,(j,a)\notin D_{i}$. Let $\alpha=(\alpha(\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))\in\mathbb{R}^{n_{0}}$ be an arbitrary vector with $\alpha||$ sufficiently small. By replacing $\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)$ and $\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)$ with $\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)$ and $\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)$ in the system (10), and subtracting $c(t)(1-\theta(t))\alpha$ from the left-hand side of the first set of equations, we derive an equivalent system formulated with fewer variables and constraints,
[…]
@@
Given almost any $\alpha\in\mathbb{R}^{n_{0}}$ with sufficiently small $\alpha||$, a smooth path can be identified in $\mathscr{P}_{L}$ that begins at $(x^{*}(2),2,\nu^{*}(2))$ when $t=2$ and leads to a normal-form perfect equilibrium as $t$ approaches zero.
@@
Let $p(x,t,\nu;\alpha)$ denote the left-hand sides of the equations in the system (16), and define $p_{\alpha}(x,t,\nu)=p(x,t,\nu;\alpha)$ when $\alpha$ is treated as a constant. The function $p(x,t,\nu;\alpha)$ is continuously differentiable over $\mathbb{R}^{n_{0}}\times(0,2)\times\mathbb{R}^{m_{0}}\times\mathbb{R}^{n_{0}}$. As demonstrated in Appendix B, its Jacobian matrix has full-row rank in this region. By using the transversality theorem as stated by Eaves and Schmedders [EavesGeneralequilibriummodels1999], one can prove that zero is a regular value of $p_{\alpha}(x,t,\nu)$ over $\mathbb{R}^{n_{0}}\times(0,2)\times\mathbb{R}^{m_{0}}$ for almost any $\alpha$ with $\alpha||<\epsilon$, where $\epsilon$ is a sufficiently small positive constant.
We select an appropriate parameter $\alpha$ such that zero is a regular value of $p_{\alpha}(x,t,\nu)$ over $\mathbb{R}^{n_{0}}\times(0,2)\times\mathbb{R}^{m_{0}}$. Applying the implicit function theorem, it follows that the component described in Theorem 5 defines a smooth path in $\mathscr{P}_{L}$, originating at $((x^{*}(2),2,\nu^{*}(2))$ and terminates at $t=0$. As demonstrated in Appendix B, the regularity of zero is preserved at $t=2$ within $\mathbb{R}^{n_{0}}\times\mathbb{R}^{m_{0}}$, thereby precluding tangential intersections with $\mathbb{R}^{n_{0}}\times\{2\}\times\mathbb{R}^{m_{0}}$. Finally, Theorem 3 confirms that this smooth path culminates in a normal-form perfect equilibrium. This completes the proof. ∎
@@ 5 Extension of Harsanyi’s tracing procedures to the sequence form
According to the analysis in [vandenelzenAlgorithmicApproachTracing1999], the tracing procedure proposed by von Stengel et al. [VonStengelComputingNormalForm2002] for two-player extensive-form games is equivalent to Harsanyi’s linear tracing procedure when applied in sequence form. In subsection 5.1, we extend this tracing procedure to accommodate $n$-player games and develop an alternative differentiable path-following method for computing normal form perfect equilibria. Furthermore, to improve the smoothness of the equilibrium paths, we extend Harsanyi’s logarithmic tracing procedure to the sequence form in subsection 5.2, providing an additional differentiable path-following method.
@@ 5.1 Harsanyi’s linear tracing procedure in sequence form
Let $p^{0}=(p^{0i}(\varpi^{i}):i\in N,\varpi^{i}\in W^{i})$ be a given prior realization plan profile. For $\gamma\in\Lambda$ and $t\in[0,1]$, define $y(\gamma,t)=(y(\gamma^{i},t;\varpi^{i}):i\in N,\varpi^{i}\in W^{i})$ with $y(\gamma^{i},t;\varpi^{i})=(1-t)\gamma^{i}(\varpi^{i})+tp^{0i}(\varpi^{i})$. This construction ensures $y(\gamma,t)\in\Lambda$. For $t\in(0,1]$, we introduce an artificial game $\Gamma_{s}^{v}(t)$ in sequence form, where each player $i$ finds their best response to $\hat{\gamma}\in\Lambda$ by solving the following linear optimization problem,
$\begin{array}[]{rl}\max\limits_{\gamma^{i}}&\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\hat{\gamma}^{-i},t))\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ &0\leq\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a),\;(j,a)\in D_{i}.\end{array}$ (17)
@@
$\begin{array}[]{l}g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\gamma^{-i},t))+\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}=0,\;i\in N,(j,a)\in D_{i},\\ g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\gamma^{-i},t))-\nu^{i}_{I^{j}_{i}}+\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,(j,a)\notin D_{i},\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)-\gamma^{i}(\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ \gamma^{i}(\varpi^{i}_{I^{j}_{i}}a)\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=0,\;0\leq\gamma^{i}(\varpi^{i}_{I^{j}_{i}}a),\;0\leq\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a),\;i\in N,(j,a)\in D_{i},\end{array}$ (18)
where $\zeta^{i}_{I^{j}_{i}}(a)=\sum_{j_{q}\in M_{i}(\varpi^{i}_{I^{j}_{i}}a)}\nu^{i}_{I^{j_{q}}_{i}}$. As a result, the solution $\gamma^{*}$ solves the optimization problem (17) against itself if and only if a pair $(\lambda^{*},\nu^{*})$ exists together with $\gamma^{*}$ that collectively satisfies the system (18). Referring to the proof of Lemma 4, we deduce that the artificial game $\Gamma_{s}^{v}(t)$ can be equivalently reformulated as the perturbed game $\Gamma_{s}(t)$ with the perturbation $\eta(t)=(\eta^{i}(t;\varpi^{i}):i\in N,\varpi^{i}\in W^{i})$ defined by $\eta^{i}(t;\varpi^{i})=tp^{0i}(\varpi^{i})$. The strategy profile $\gamma^{*}$ is a Nash equilibrium of $\Gamma_{s}^{v}(t)$ whenever $y(\gamma^{*},t)$ constitutes a Nash equilibrium in $\Gamma_{s}(t)$. This leads to the following theorem.
@@
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(t)}&(1-\theta(t))\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\hat{\gamma}^{-i},\rho(t)))\\ &-\theta(t)\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)\ln\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ &0\leq\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a),\;(j,a)\in D_{i}.\end{array}$ (19)
@@
Applying the optimality conditions to the problem (19), together with $\hat{\gamma}=\gamma(t)$, yields a polynomial equilibrium system. This system is specified by (20) for $t\in(1,2]$ and is equivalent to (18) for $t\in(0,1]$.
@@
$\begin{array}[]{l}(1-\theta(t))g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\gamma^{-i}(t),\rho(t)))+\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)\\ \hskip 105.2751pt-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i},\\ \gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a),\;0<\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (20)